# FOSP: Fine-tuning Offline Safe Policy through World Models

### v1

@@ 4.1 In-sample Optimization for Offline Training
where ${\bm{r}}_{t}$ is the reward from the offline dataset. Finally, the critic loss is the combination of the above.
$\mathcal{L}_{Q}(\phi)=\mathcal{L}_{Q}^{1}(\phi)+\mathcal{L}_{Q}^{2}(\phi)$ (8)
4.2 Model-based Policy Expansion for World Models Fine-tuning
After offline pretrain, we fine-tune the world model with online exploration to enhance performance and improve its generalization to similar tasks. However, directly using online algorithms suffers from non-recoverable performance drops potentially caused by the distribution shift and the change of the dynamic model. On the other hand, standard offline RL algorithms typically exhibit worse online fine-tuning performance due to their conservative design[20, 17]. Therefore, we try to bridge offline-online RL by fully utilizing offline pretraining policies[32].
Suppose that $\pi_{\beta}$ is the offline policy. Instead of directly fine-tuning it in the online stage, we optimize a new policy $\pi_{\psi}$ by gradually learning from prior policy $\pi_{\beta}$ and online exploring simultaneously. Concretely, given a current state ${\bm{s}}$, an action set is composed of actions generated by two policies: $\mathcal{A}=\{{\bm{a}}_{\beta}\sim\pi_{\beta}({\bm{s}}),{\bm{a}}_{\psi}\sim\pi_{\psi}({\bm{s}})\}.$ The actions are selected probabilistically based on their potential utilization value:
$P(\omega)[k]=\frac{\exp(Q_{\phi}({\bm{s}},{\bm{a}}_{k})/\alpha)}{\exp(Q_{\phi}({\bm{s}},{\bm{a}}_{\beta})/\alpha)+\exp(Q_{\phi}({\bm{s}},{\bm{a}}_{\psi})/\alpha)},\;\forall k\in\{\beta,\psi\},$ (9)
where $\alpha$ is temperature. According to this probability function, the two policies can be trained in a joint adaptive manner that has the advantage of adaptiveness in the sense of allowing generalizing of new abilities. In this way, the offline replay-buffer can be directly transferred to the online fine-tuning stage simplifying the complexity of balanced sampling[26].
In the model-based setting, the dynamic model also needs to be fine-tuned since the model primarily generates the policy update data. The world model benefits from its robust design in capturing temporal dependencies and modeling latent dynamics, which contributes to its stability during fine-tuning. However, the dynamic model might be influenced by the incorrect estimates learned from conservative offline training, making online fine-tuning more challenging. Thus, we develop an additional dynamic model for fine-tuning, expanding the policy to adapt to new tasks.
@@ 4.3 Reachability Estimation Function as Safety Guarantee
To ensure safety throughout the entire training process, prior works consider adopting the relaxation of the Lagrangian multiplier method on (4). The optimization problem is transformed into an unconstrained safe model-based problem with the Lagrangian multiplier $\lambda_{i}$:
@@
However, the soft constraints will still lead to a certain chance of violation. It becomes more severe in the offline setting because of the difficulty in estimating the cost values. Hence, we replace soft constraints with hard constraints by the feasibility guidance approach. We first divide our optimization problem into two parts: the feasible part and the infeasible part.
For the feasible part, we want the agent to maintain safety within the feasible set and maximize the reward return. For the infeasible part, we want the policy to avoid the violations as little as possible and return to the feasible set. As the safe offline RL problem has another objective to optimize: $D(\pi_{\psi}||\pi_{b})$ ($\pi_{b}$ is the behavior policy), we can reformulate our offline safe model-based problem as follows:
@@
where $V_{\varphi}^{r}({\bm{s}})=\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi_{\psi}}[\sum_{t=0}^{\infty}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})]$ and $V_{\varphi}^{c}({\bm{s}})=\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi_{\psi}}[\sum_{t=0}^{\infty}\gamma^{t}c({\bm{s}}_{t},{\bm{a}}_{t})]$ represent the expected cumulative rewards and costs, $\mathcal{S}_{f}$ is the feasible set containing all feasible states. Note that the policy constraint is a soft constraint so the feasible part becomes a problem incorporating coupled hard and soft constraints. Thus, we use the reachability estimation function $u^{\pi}({\bm{s}})=\operatorname*{\mathbb{E}}_{\tau\sim\pi}[\max_{{\bm{s}}_{t}\in\tau}\mathds{1}_{({\bm{s}}_{t}|{\bm{s}}_{0},\pi)\in\mathcal{S}_{f}}]$ to handle the mixed constraints while prioritizing the hard constraints from RESPO[51]. It also assists the agent in reentering the feasible region even when it is in the infeasible region.
Therefore, we couple the out-of-distribution actions avoidance into the problem by using advantage functions. Also, we take advantage of the Augmented Lagrangian with a proximal relaxation method that is quite effective in safe model-based RL algorithms. The final problem is as follows:
$\max_{\pi_{\psi}}\;\operatorname*{\mathbb{E}}_{{\bm{s}}}\Big{[}(A^{r}({\bm{s}},{\bm{a}})-\Phi(V_{\varphi}^{c}({\bm{s}}),\lambda_{p}^{k},\mu^{k}))\cdot u^{\pi}_{\psi}-A^{c}({\bm{s}},{\bm{a}})\cdot(1-u^{\pi}_{\psi})\Big{]},\;\;\text{s.t.}\;\;D(\pi_{\psi}||\pi_{b})\leq\epsilon.$ (13)
where $A^{r}({\bm{s}},{\bm{a}})=Q_{\phi}^{r}({\bm{s}},{\bm{a}})-V_{\varphi}^{r}({\bm{s}})$ and $A^{c}({\bm{s}},{\bm{a}})=Q_{\phi}^{c}({\bm{s}},{\bm{a}})-V_{\varphi}^{c}({\bm{s}})$. We can finally obtain a closed-form solution of $\pi_{\psi}$, which we can use to extract the optimal policy. (For more details see Appendix.A.1)
@@
$\mathcal{L}_{\pi}(\psi)=-\operatorname*{\mathbb{E}}_{\tau\sim\mathcal{D}}\Big{[}\sum_{t=1}^{T}(w\cdot\log\pi_{\psi}({\bm{a}}_{t}|{\bm{s}}_{t})+\eta H[\pi_{\psi}({\bm{a}}_{t}|{\bm{s}}_{t})])-\Phi(V_{\varphi}^{c}({\bm{s}}),\lambda_{p}^{k},\mu^{k})\cdot u^{\pi}_{\psi}\Big{]},$ (15)
@@
In the online fine-tuning stage, we apply policy expansion as follows: use $\pi_{b}$ as the frozen offline policy trained by (15) and learn a new policy $\pi_{\psi}$ without policy behavior constraint in (13). Thus, the online policy loss can change the advantage weight as:
@@
For cost critic $Q^{c}(s,a)$ and value $V^{c}(s)$ training, we use the same way as Section 4.1. The whole algorithm training outline is shown in Appendix.B.
Figure 2: Offline experimental results. Comparing FOSP to baselines across five image-based safety tasks. The results for all three algorithms are obtained after training for 1 million steps. Reward: averaged episode reward return. Cost: averaged episode cost return. Cost Regret: averaged cost value throughout the training phase.
@@ 5 Experimental Results
In this section, we evaluate our method on different agents and tasks from the Safety-Gymnasium in simulation. We deploy it in motion planning control tasks on a real Franka robot as well. We aim to address the following issues: (1) How does FOSP perform in offline and online fine-tuning? (2) How does each component in FOSP contribute to its overall functionality? (3) Does FOSP successfully handle unseen safety regions in the real world?
@@ 5.1 Experimental Setup
We consider five tasks on Safety-Gymnasium benchmark[11] environments. The agents need to navigate to predetermined goals without collision with hazards and vases. We measure the performance with three metrics: average undiscounted episode reward return (Reward), average undiscounted episode cost return (Cost) and average cost value throughout the training phase (Cost Regret). To evaluate the advantage of world models in handling high-dimensional features, we use 64$\times$64 pixels RGB image as inputs obtained from the first-person perspective of the agent, which is more representative of real-world environments. The offline datasets are sampled by three different behaviors: unsafe policy, safe policy and random policy to evaluate the performance of FOSP on diverse behavior samples. They are mixed in a 1:1:1 ratio, with each part containing 200 trajectories. For more information see Appendix.C.
@@
Because our work is the first approach in the safe model-based offline-to-online area to the best of our knowledge, there are not many baselines for us to compare. Prior works in offline safe RL use the model-free method which typically performs poorly in vision-only tasks. Therefore, we compare FOSP with the following two model-based baselines: SafeDreamer[8], an online powerful algorithm in safe model-based reinforcement learning without online planning and DreamerV3[5], a strong baseline in visual tasks but overlooks constraint violations.
Figure 3: Ablation studies. We evaluate ablations in SafetyPointGoal2 with means of three seeds. The vertical line divides the offline and online phases. (1M steps for offline and 0.75M steps for online)
@@
The real robot environment includes a 7-DOF Franka Emika Panda robot as the controlled agent and a static third-person Intel RealSense camera to obtain images as inputs. Also, the agent has additional inputs: the robot end-effector pose information. We design the task SafeReach that controls the robot to reach a desired goal while avoiding collisions with predefined obstacles in its field of view. The rewards and costs are sparse and detected manually determining whether the robot reaches the goal or violates the constraints. The dataset consists of demonstration data, safety violation data, and failure data. Furthermore, we design some unseen safe regions and different targets that are not contained in the offline dataset. These unseen tasks are related to some undefined obstacles that need to be identified through observation which can not be solved by motion planning methods. The task further illustrates the practicality of our approach.

### v2

@@ 4.1 In-sample Optimization for Offline Training
where ${\bm{a}}_{t},{\bm{r}}_{t}$ are from the offline dataset and the final critic loss becomes
$\mathcal{L}_{Q}(\phi)=\mathcal{L}_{Q}^{1}(\phi)+\mathcal{L}_{Q}^{2}(\phi),$ (11)
@@ 4.2 Reachability Estimation Function as Safety Guarantee
To ensure the safety of the policy, prior works consider adopting the relaxation of the Lagrangian multiplier method (Stooke et al., 2020). The optimization problem on equation 4 can be transformed into an unconstrained problem with the Lagrangian multiplier $\lambda_{i}$:
@@
However, the soft constraints will still lead to a certain chance of violation, which is exacerbated in the offline setting due to the difficulty in estimating the cost values. Hence, we replace soft constraints with hard constraints by the feasibility guidance approach (Yu et al., 2022). The feasible states are included in the feasible set $\mathcal{S}_{f}({\bm{s}}):=\{{\bm{s}}|V_{\varphi}^{c}({\bm{s}})=0\}$. Then we can divide the optimization problem into two parts: the feasible part and the infeasible part.
For the feasible part, we want the agent to maintain safety within the feasible set and maximize the reward return. For the infeasible part, we want the policy to violate the constraints as little as possible and return to the feasible set. As the safe offline RL problem has another objective to optimize: $D_{\mathrm{KL}}(\pi_{\psi}||\pi_{b})$ ($\pi_{b}$ is the behavior policy), we can reformulate the problem as follows:
@@
Note that the policy constraint $D_{\mathrm{KL}}(\pi_{\psi}||\pi_{b})\leq\epsilon$ is a soft constraint, which complicates the feasible part problem due to coupled hard and soft constraints. Thus, we use the reachability estimation function $u^{\pi}({\bm{s}})=\mathbb{E}_{\tau\sim\pi}[1-\max_{{\bm{s}}_{t}\in\tau}(\mathds{1}\{({\bm{s}}_{t}|{\bm{s}}_{0},\pi)\in\mathcal{S}_{v}\})]$ from RESPO (Ganai et al., 2024), where $\mathcal{S}_{v}$ and $\mathcal{S}_{f}$ are complementary. This function helps handle the mixed constraints by prioritizing the hard constraints and also assists the agent in reentering the feasible region even when it ventures into the infeasible region. And the problem becomes:
$\max_{\pi_{\psi}}\;\mathbb{E}_{{\bm{s}}}\Big{[}V_{\varphi}^{r}({\bm{s}})\cdot u^{\pi}_{\psi}-V_{\varphi}^{c}({\bm{s}})\cdot(1-u^{\pi}_{\psi})\Big{]},\;\;\text{s.t.}\;\;V_{\varphi}^{c}({\bm{s}})=0,\;\;D_{\mathrm{KL}}(\pi_{\psi}||\pi_{b})\leq\epsilon.$ (15)
To simplify the constraints, we introduce Augmented Lagrangian with a proximal relaxation method. Meanwhile, we leverage advantage functions $A^{r}({\bm{s}},{\bm{a}})=Q_{\phi}^{r}({\bm{s}},{\bm{a}})-V_{\varphi}^{r}({\bm{s}})$ and $A^{c}({\bm{s}},{\bm{a}})=Q_{\phi}^{c}({\bm{s}},{\bm{a}})-V_{\varphi}^{c}({\bm{s}})$ to assess the influence of actions on the value function according to Proposition 1. The final problem can be formulated as follows:
$\max_{\pi_{\psi}}\;\mathbb{E}_{{\bm{s}}}\Big{[}(A^{r}({\bm{s}},{\bm{a}})-\Phi(V_{\varphi}^{c}({\bm{s}}),\lambda_{p}^{k},\mu^{k}))\cdot u^{\pi}_{\psi}-A^{c}({\bm{s}},{\bm{a}})\cdot(1-u^{\pi}_{\psi})\Big{]},\;\;\text{s.t.}\;\;D_{\mathrm{KL}}(\pi_{\psi}||\pi_{b})\leq\epsilon.$ (16)
$\Phi\left(V_{\varphi}^{c},\lambda_{p}^{k},\mu^{k}\right)=\begin{cases}\lambda_{p}^{k}V_{\varphi}^{c}+\frac{\mu^{k}}{4}(V_{\varphi}^{c})^{2}&\text{ if }\lambda_{p}^{k}+\frac{\mu^{k}}{2}V_{\varphi}^{c}\geq 0,\\ -\frac{\left(\lambda_{p}^{k}\right)^{2}}{\mu^{k}}&\text{ otherwise }.\end{cases}$ (17)
where $\lambda_{p}^{k}$ and $\mu^{k}$ are Lagrange multipliers. We can finally obtain a closed-form solution of $\pi_{\psi}$, which we can use to extract the optimal policy (Zheng et al., 2023). (For more details see Appendix.A.2)
@@
and the policy loss is as follows:
$\mathcal{L}_{\pi}(\psi)=-\mathbb{E}_{\tau\sim\mathcal{D}}\Big{[}\sum_{t=1}^{T}(w\cdot\log\pi_{\psi}({\bm{a}}_{t}|{\bm{s}}_{t})+\eta E[\pi_{\psi}({\bm{a}}_{t}|{\bm{s}}_{t})])-\Phi(V_{\varphi}^{c}({\bm{s}}),\lambda_{p}^{k},\mu^{k})\cdot u^{\pi}_{\psi}\Big{]},$ (20)
where $\eta$, $\beta_{1}$, $\beta_{2}$ are hyperparameters, $E$ is the entropy. For critic learning, we use the same way to train cost critic $Q^{c}_{\phi^{\prime}}({\bm{s}},{\bm{a}})$ and cost value $V^{c}_{\varphi^{\prime}}({\bm{s}})$ as Section 4.1.
4.3 Safe Policy Expansion for World Models Fine-tuning
We aim to improve performance and generalize the learned policy to similar tasks while maintaining the safety learned from the offline policy through online fine-tuning. However, applying online safe RL algorithms for fine-tuning often leads to initial performance drops and constraint violations, probably caused by the distribution shift between the offline and online data (Lee et al., 2022b). Conversely, directly fine-tuning the original offline RL algorithm typically exhibits worse online performance due to their conservative design (Kostrikov et al., 2021; Kumar et al., 2020). Therefore, we try to bridge offline-online RL by safe policy expansion.
Suppose that $\pi_{\psi}$ is the offline learned policy. Rather than directly fine-tuning it in the online stage, we optimize a new policy $\pi_{\psi^{\prime}}$ by gradually learning from frozen prior policy $\pi_{\psi}$. Concretely, given a current state ${\bm{s}}$, an action set from which actions are selected is composed of actions generated by two policies: $\mathcal{A}=\{{\bm{a}}_{\psi}\sim\pi_{\psi}({\bm{s}}),{\bm{a}}_{\psi^{\prime}}\sim\pi_{\psi^{\prime}}({\bm{s}})\}.$ The actions are then probabilistically selected based on their potential utilization value (Zhang et al., 2023):
$P(\omega)[k]=\frac{\exp(\tilde{Q}({\bm{s}},{\bm{a}}_{k})/\alpha)}{\exp(\tilde{Q}({\bm{s}},{\bm{a}}_{\psi})/\alpha)+\exp(\tilde{Q}({\bm{s}},{\bm{a}}_{\psi^{\prime}})/\alpha)},\;\forall k\in\{\psi,{\psi^{\prime}}\},$ (21)
$\tilde{Q}({\bm{s}},{\bm{a}})=\frac{Q_{\phi}({\bm{s}},{\bm{a}})}{\hat{Q}^{*}}-\frac{Q^{c}_{\phi^{\prime}}({\bm{s}},{\bm{a}})}{\hat{C}^{*}},$ (22)
where $\alpha$ is temperature. We use upper bounds $\hat{Q}^{*}=\sum_{t=0}^{T}\gamma^{t}r_{max}$, $\hat{C}^{*}=\sum_{t=0}^{T}\gamma^{t}c_{max}$ to normalize the Q-value, where $r_{max},c_{max}$ represent the max reward and cost. $\tilde{Q}({\bm{s}},{\bm{a}})$ represents the balance between the normalized cumulative rewards and cumulative costs. Based on the probability function, the two policies can adaptively explore the environment safely according to their respective probabilities, determined by the Q-values of their generated actions. We can ensure the stability of initial performance during the early fine-tuning stages, exploring with a better policy under the premise of minimal constraint violations. At the same time, we leverage online algorithms to enhance performance further. According to Section 4.2, the weight of policy loss in the online stage derived from equation 16 without the behavior policy constraint can be written as:
@@ 4.3 Safe Policy Expansion for World Models Fine-tuning
This approach learns a new policy through online updates while leveraging the offline-trained policy as prior knowledge. Ultimately, this enables the policy to generalize to new tasks during online fine-tuning safely. In this way, the offline replay buffer can be directly converted into an online replay buffer, allowing online exploration trajectories to be continuously added during online fine-tuning, thereby simplifying the complexity of balanced sampling (Lee et al., 2022b).
@@ 5 Experimental Results
In this section, we evaluate our method across different agents and tasks within the Safety-Gymnasium simulation environment (Ji et al., 2023), as well as on real-world motion planning control tasks using a Franka robot. We aim to address the following issues: (1) How does FOSP perform in offline pretrain and online fine-tuning? (2) What role does each component of FOSP play in its overall functionality? (3) How does the offline dataset affect experimental results? (4) Can FOSP successfully handle unseen safety regions in the new scenarios?
@@ 5.1 Experimental Setup
Table 1: Offline-to-online results. Reward return and cost return of methods in offline-to-online fine-tuning (1M steps for offline and 0.5M steps for online fine-tuning). We report the mean value of 5 independent runs with different seeds.
FOSP(Ours) Recover-RL DreamerV3 SafeDreamer
Tasks Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$
PointGoal1 18.7$\rightarrow$21.5 1.4$\rightarrow$0.2 5.4$\rightarrow$12.6 1.8$\rightarrow$0.8 17.4$\rightarrow$25.8 82.3$\rightarrow$85.1 12.7$\rightarrow$19.2 1.5$\rightarrow$0.02
PointButton1 14.7$\rightarrow$18.1 9.5$\rightarrow$2.1 5.4$\rightarrow$6.3 10.3$\rightarrow$2.8 15.1$\rightarrow$20.8 167.7$\rightarrow$157.9 8.9$\rightarrow$10.7 10.2$\rightarrow$4.5
PointPush1 4.0$\rightarrow$13.2 18.1$\rightarrow$0.1 0.4$\rightarrow$3.8 24.8$\rightarrow$0.5 2.3$\rightarrow$14.6 30.9$\rightarrow$26.3 1.3$\rightarrow$10.3 17.8$\rightarrow$0.12
PointGoal2 8.1$\rightarrow$13.5 7.5$\rightarrow$0.23 0.5$\rightarrow$2.7 53.2$\rightarrow$0.3 9.5$\rightarrow$18.9 367.1$\rightarrow$290.2 3.6$\rightarrow$12.7 8.6$\rightarrow$0.3
CarGoal2 10.1$\rightarrow$14.5 1.6$\rightarrow$0.07 4.6$\rightarrow$8.9 0.6$\rightarrow$0.13 16.7$\rightarrow$24.1 231.0$\rightarrow$287.9 6.9$\rightarrow$9.8 0.7$\rightarrow$0.09
Average 16.6 0.54 6.7 0.91 20.8 171.1 12.5 1.01
Table 2: Compare with online algorithms. We compare our method with some online-only algorithms. SafeDreamer(+planning) was trained online for 0.5M steps and CPO and PPO-Lagrangian were trained for 10M training steps until they converged. We report the mean value of 5 independent runs with different seeds.
FOSP(Ours) SafeDreamer(+planning) CPO PPO-Lag
Tasks Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$ Reward $\uparrow$ Cost $\downarrow$
PointGoal1 21.5 0.2 20.1 0.6 22.3 45 19.5 29
PointButton1 18.1 2.1 12.4 5.5 17.1 76 5.3 25
PointPush1 13.2 0.1 8.1 0.7 3.7 35 2.9 26
PointGoal2 13.5 0.23 13.0 1.7 13.8 51 1.8 30
CarGoal2 14.5 0.07 11.2 0.34 15.5 52 7.8 28
Average 16.6 0.54 12.9 1.77 14.5 51.8 7.5 27.7
Figure 3: Online experimental results. Comparing FOSP to baselines across five image-based safety tasks at the online fine-tuning stage. The results for model-based algorithms are obtained after fine-tuning for 750,000 steps. The dashed lines represent the benchmark results for CPO and PPO-Lagrangian after 10 million training steps across all tasks. The SafeDreamer (planning) was trained online for 0.75 million steps. Reward: averaged episode reward return. Cost: averaged episode cost return. Cost Regret: averaged cost value throughout the training phase.
@@
We consider five tasks on Safety-Gymnasium benchmark (Ji et al., 2023) environments. The agents need to navigate to predetermined goals without collision with hazards and vases. We measure the performance with three metrics: average undiscounted episode reward return (Reward), average undiscounted episode cost return (Cost) and average cost value throughout the training phase (Cost Regret). To evaluate the advantage of world models in handling high-dimensional features, we use 64$\times$64 pixels RGB image as inputs obtained from the first-person perspective of the agent, which is more representative of real-world environments. The standard offline datasets are sampled by three different behaviors: unsafe policy, safe policy and random policy. They are mixed in a 1:1:1 ratio, with each part containing 200 trajectories. All the tasks are trained for five seeds. For more information see Appendix.D.
Figure 4: Module ablation studies. We evaluate ablations in SafetyPointGoal2 with means of five seeds. The vertical line divides the offline and online phases.
@@
Prior works in offline safe RL using the model-free method typically perform poorly in vision-only tasks due to their slow convergence rate (Liu et al., 2023; Zheng et al., 2023). Therefore, we mainly compare FOSP with the following model-based baselines: Recovery RL (model-based version) (Thananjeyan et al., 2021), a method first learns the constraint violation regions offline and then performs online policy training; SafeDreamer (Huang et al., 2023), an online powerful algorithm in safe model-based reinforcement learning and DreamerV3 (Hafner et al., 2023), a strong baseline in visual tasks but overlooks constraint violations. We also choose classical safe RL algorithms CPO (Achiam et al., 2017) and PPO-Lagrangian (Schulman et al., 2017) to compare at the online stage, following the experimental protocol of Ray et al. (2019).
@@
The real robot environment includes a 7-DOF Franka Emika Panda robot as the controlled agent and a static third-person Intel RealSense camera to obtain images as inputs. The agent also receives additional inputs, specifically the pose information of the robot’s end-effector. We design the task SafeReach that controls the robot to reach a desired goal while avoiding collisions with predefined obstacles in its field of view. The rewards and costs are sparse and detected manually, determining whether the robot reaches the goal or violates the constraints. The dataset consists of demonstration data, violation data, and failure data. Furthermore, we design some unseen safe regions and different targets that are not contained in the offline dataset. These unseen tasks require the agent to recognize objects that it has not encountered in the dataset and to generalize its decision-making abilities to these new scenarios. The tasks further illustrate the practicality of our approach.